Circular No. 6606

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1750.8–2900 AND GRS 1737–31 J. Heise, on behalf of the BeppoSAX Wide Field Camera (WFC) team

(composed of individuals at the Space Research Organization of the Netherlands, Utrecht; at Istituto di Astrofisica Spaziale, CNR, Frascati; and at BeppoSAX, Rome) reports on observations made with the BeppoSAX WFC between Mar. 19.197 and 20.775 UT: “An improved position of the soft transient x-ray source SAX J1750.8–2900 (IAUC 6597) is α = 17^{h}50^{m}24^{s}, δ = −29^{o}01′.9 (equinox 2000.0; error radius 3′), which is 6′.2 from the variable VLA radio source reported on IAUC 6601, and 2′.3 from 1RXP J175029– 2859.9. Average intensities were 26 and 22 mCrab at 1.6–7.7 and 7.7–23 keV, respectively. The hard x-ray transient source GRS 1737–31 (IAUC 6599) was also observed with the WFC, yielding an average intensity of 11 mCrab over 1.6–7.7 keV and 22 mCrab over 7.7–23 keV. The source is located at α = 17^{h}40^{m}00^{s}, δ = −31^{o}04′.2 (equinox 2000.0; error radius 4′), which is 5′.6 from the reported position on IAUC 6599.”

GRB 970228 K. Sahu, M. Livio, L. Petro, and F. D. Macchetto, Space Telescope

Science Institute, write: “GRB 970228 was observed with the HST Wide Field and Planetary Camera (WFPC2), between Mar. 26.11 and 26.28 UT, ∼ 26 days after the outburst. The optical transient (α = 5^{h}01^{m}46^{s}.7, δ = +11^{o}46′53′′.0, equinox 2000.0; cf. IAUC 6584, 6588) was placed at the center of the Planetary Camera field-of-view, so as to attain maximum angular resolution. The other three candidates (two optical and one radio) proposed earlier as counterparts of GRB 970228 (IAUC 6576, 6574, 6577) are also within the WFPC2 field-of-view. Observations were taken with the F606W (wide V) and F814W (I) filters. Four exposures with a total integration time of 4700 s were taken through the F606W filter and two exposures with a total integration time of 2400 s were taken through the F814W filter. The optical counterpart is clearly detected in both passbands. It was found to be embedded in an extended source, which can also be seen. The derived magnitudes for the point source are 25.7 ± 0.3 and 24.2 ± 0.3 in the F606W and F814W filters, respectively. The data were put into the HST archive on the day of the observations. The data can be accessed

through the Web at http://www.stsci.edu/top.html.”

1997 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green